OpenStudy (anonymous):

Solve ln(5x + 7) = 8. Round to the nearest thousandth.

OpenStudy (blurbendy):

Cancel logs by taking exp of both sides 5x + 7 = e^8 Subtract 7 from both sides 5x = e^8 -7 Solve for x x = e^8 / 5 - 7/5

OpenStudy (blurbendy):

x = 594.792

OpenStudy (anonymous):

alright thank you for your help!

OpenStudy (blurbendy):

No probs.